# Sampling 3D Molecular Conformers with Diffusion Transformers

### v1

@@ 4.2 Positional Embeddings ( PE )
where $\vec{r}_{i}\in\mathbb{R}^{3}$ is the Cartesian position of the $i$-th atom. This kind of PE is neither rotationally nor translationally invariant, and serves as a baseline without any symmetry constraints.
Relative Positional Embedding ( rPE )
We use displacements vectors $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ to build pairwise relative PEs as
$\boldsymbol{p}_{ij}^{\text{rPE}}=\text{MLP}(\vec{r}_{ij})\,.$ (12)
This formulation ensures translational invariance but not rotational invariance.
Euclidean Positional Embedding ( PE(3) )
Adapting ideas from equivariant message passing neural networks like PaiNN [61] or NequIP [62], we construct SO(3)-equivariant pairwise PEs as a concatenation of $L+1$ components
@@
where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0\dots L$. The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$ and (after concatenation) $\boldsymbol{p}_{ij}^{\text{PE(3)}}\in\mathbb{R}^{(L+1)^{2}\times H}$. Under rotation of the input positions, these PEs transform equivariantly (see Appendix B.2.2). Moreover, because displacement vectors are used as inputs, the embeddings are also invariant to translations. As a result, they respect the full set of Euclidean symmetries relevant to molecular geometry.
4.3 Self-Attention Operation
For ease of notation, we only describe self-attention with a single head, but employ multi-head attention [29] with $n_{\text{heads}}$ heads in our experiments. All self-attention blocks rely on query, key and value vectors, which are obtained from the input tokens =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\ |\ \boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ as
$\boldsymbol{q}=\boldsymbol{W}_{q}\boldsymbol{\tilde{h}}\,,\hskip 25.0pt\boldsymbol{k}=\boldsymbol{W}_{k}\boldsymbol{\tilde{h}}\,,\hskip 25.0pt\boldsymbol{v}=\boldsymbol{W}_{v}\boldsymbol{\tilde{h}}\,,$ (14)
where $\boldsymbol{W}_{q},\boldsymbol{W}_{k},\boldsymbol{W}_{v}\in\mathbb{R}^{H\times H}$ are trainable weight matrices and $\boldsymbol{\tilde{h}}$ is either identical to $\boldsymbol{h}$ or combines it with a PE (see below). We define a slightly modified similarity kernel
$\text{sim}(\boldsymbol{q},\boldsymbol{k},\boldsymbol{u})=\exp\Bigg{(}\frac{\boldsymbol{q}^{\intercal}\cdot(\boldsymbol{k}\,\odot\boldsymbol{u})}{\sqrt{H}}\Bigg{)},$ (15)
where $\boldsymbol{u}\in\mathbb{R}^{H}$ is used to inject additional information, e.g., conditioning signals and/or positional embeddings, and $\odot$’ denotes element-wise multiplication. Depending on the subset of Euclidean symmetries we aim to incorporate, we adopt slightly different formulations of the self-attention mechanism, as detailed below.
Standard Self-Attention
For absolute and relative PEs, we slightly modify standard self-attention to allow injecting pair-wise information into the values in addition to using our modified similarity kernel:
$\text{ATT}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\boldsymbol{v}_{j}\odot\boldsymbol{u}_{ij})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,.$ (16)
Queries, keys, and values are obtained with Eq. 14 from different (position-encoded) tokens $\boldsymbol{\tilde{h}}_{i}$ depending on the chosen PEs; further, the injected pair-wise information $\boldsymbol{u}_{ij}$ differs:
[…]
Here $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ are pair-wise graph conditioning tokens (see Eqs. 9 and 10) and $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$ and $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$ are the absolute and relative PEs described above (see Eqs. 11 and 12).
SO(3)-Equivariant Self-Attention
To preserve all Euclidean symmetries throughout the network, every token must transform equivariantly. One way to achieve this is by separating out the rotational degrees of freedom, encoding them with irreducible representations of the rotation group $\mathrm{SO(3)}$. This introduces a “degree-axis” of size $(L+1)^{2}$, which encodes angular components of increasing order. The maximum degree $L$ is chosen to ensure high fidelity at a reasonable computational cost. For example, setting $L=1$ restricts the representation to scalars and vectors, as used in models like PaiNN [61] or TorchMDNet [34]. An $\mathrm{SO(3)}$-equivariant formulation of self-attention is then given as
$\text{ATT}_{\text{SO(3)}}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\hat{\boldsymbol{u}}_{ij}\otimes\boldsymbol{v}_{j})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,,$ (18)
where equivariant queries, keys and values can be calculated similarly to Eq. 14 and $\otimes$’ denotes a Clebsch-Gordan (CG) tensor product contraction [60]. The dot-product in the similarity measure is taken along both feature and degree axes, such that the overall update preserves equivariance (see Appendix for details). Tokens and scaling vectors are calculated as
[…]
where $\boldsymbol{\phi}(r_{ij})\in\mathbb{R}^{(L+1)^{2}\times H}$ is a radial filter, and the element-wise products with the pair-wise conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{1\times H}$ are broadcast along the degree axis. Importantly, the $12\ell+$ 1 subcomponents of the radial filter for degree $\ell$ are obtained by repeating per-degree filter functions $\boldsymbol{\phi}_{\ell}(r_{ij})\in\mathbb{R}^{1\times H}$ along the degree axis to preserve equivariance (see also Eq. 13).
5 Experiments
Datasets and Metrics
We conduct our experiments on the GEOM dataset [63], comprising QM9 (133,258 small molecules) and AICures (304,466 drug-like molecules). Reference conformers are generated using CREST [64]. Drug-like molecules exhibit greater structural diversity, including more rotatable bonds and multiple stereocenters. Data splits are taken from Ref. [65].
We evaluate our models’ ability to generate accurate and diverse conformers using Average Minimum RMSD (AMR) and Coverage (COV), measuring Recall (ground-truth coverage) and Precision (generation accuracy). A generated conformer is considered valid if it falls within a specified RMSD threshold of any reference conformer ($\delta=0.5\mbox{\AA}$ for GEOM-QM9 and $\delta=0.75\mbox{\AA}$ for GEOM-DRUGS). Following prior work, we generate $2K$ conformers per test molecule with $K$ reference structures. Appendix D.3 provides further details on the calculation of metrics. Following ET-Flow [32] and GeomMol [65] we also apply chirality correction (see Appendix D.4).
Conditioning Strategies
To assess the impact of different graph conditioning strategies, we compare “node only”, “node & bond-pair”, and “node & all-pair” conditioning (see Sec. 4.1 for details) against a baseline without graph conditioning. We find all variants to be effective, with “node & all-pair” conditioning performing best (see Appendix Tab. A8). It outperforms both, “node only” and “node & bond-pair” conditioning, highlighting the strength of our all-pair-conditioning strategy based on graph geodesics. This underlines the importance of deriving conditioning tokens for all atom pairs, not just those connected by edges in the molecular graph, which lack global information about the graph structure.
@@ 5 Experiments
Table 1: Results on GEOM-QM9 for different generative models (parameter counts in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Tab. A9 for results including standard deviations.
@@
Figure 2: Analysis of equivariant (PE(3)) and non-equivariant (aPE, rPE) model formulations. (A) Mean coverage recall (COV-R) versus threshold $\delta$ (root mean square deviation (RMSD) to any reference conformer). (B) Histogram of the minimal RMSD per generated sample. (C) Loss as a function of latent time $t$ relative to PE(3) loss.
Ablating Self-Attention and Positional Embedding Strategies
@@
The modular structure of DiTMC allows efficient exploration of the design space through variations in the PE strategy and associated attention blocks (see Sec. 4). All DiTMC models share hyperparameters, differing only in the choice of PEs and self-attention formulation. For the architecture using PE(3), we reduce the number of heads to match the parameter count across models. Tab. 1 shows results on GEOM-QM9. All DiTMC models produce diverse, high quality samples, outperforming the current SOTA in all AMR metrics and COV-P. This demonstrates the broad applicability of our modular design and conditioning. We use the harmonic prior introduced in Ref. [66] throughout, which yields improved results (see Appendix Tab. A15). Therefore, we use it in all following experiments.
Probing the effect of Euclidean symmetries
@@
Our PEs form a hierarchy based on the extent of Euclidean symmetry incorporated by construction. This enables a systematic evaluation of how incorporating symmetry affects model behavior. We summarize our findings below.
@@
Equivariance improves the fidelity of samples. We compute COV-R as a function of threshold $\delta$ for different PEs and self-attention blocks (Fig. 2A). SO(3)-equivariant attention with PE(3) outperforms the other variants at low $\delta$, indicating that many of the generated geometries closely match the ground-truth structures. This appears as a leftward shift in the distribution of the minimal RMSD found per generated structure (Fig. 2B) and aligns with the observation that the PE(3) model gives better AMR-R and AMR-P values (see Tab. 1). To better understand this behavior, we analyze the loss over time $t$ and find that models with non-equivariant PEs exhibit higher error near the data distribution ($t=1$) (Fig. 2C). The increase in error towards the end of the generation trajectory results in noisier structures and reduced fidelity.
@@
Equivariance increases the computational cost for models of similar size. The benefit of higher fidelity comes at increased computational cost. During training, the equivariant PE(3) model is approximately $3.5$ times slower than models using aPE or rPE, while at inference time, the factor is $\sim 3$ (see Appendix Tab. A5). All models use the same number of layers and differ only in the number of heads per layer in order to match the total parameter count.
@@
Table 2: Results on GEOM-DRUGS for different generative models (parameter counts in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Tab. A10 for results including standard deviations.
@@
*DiTMC+aPE-L 79.2 84.4 0.432 0.386 77.8 85.7 0.470 0.387
*DiTMC+PE(3)-L 80.8 85.6 0.415 0.376 76.4 82.6 0.491 0.414
Figure 3: (A) Time per gradient step for the DiTMC S, M and L variants using aPE and PE(3), measured on an Nvidia H100 GPU (batch size 128 on Geom QM9). (B) Coverage precision and recall as function of threshold $\delta$ and (C) mean AMR metrics for base (“B”) and large (“L”) model of DiTMC+aPE.
Drug-like Molecules
Finally, we evaluate our architecture on the challenging GEOM-DRUGS dataset. We analyze its behavior when scaling model capacity, assess physical validity via ensemble property prediction, and examine out-of-distribution performance. We focus on the models using aPE and PE(3), as they represent the extreme cases of not including any, and including all Euclidean symmetries in the architecture design, respectively.
Model Scaling. To analyze how DiTMC behaves as model capacity scales, we define a small base (“B”) and a large (“L”) model variant (see Appendix B). The compute cost per gradient step for the equivariant PE(3) model is substantially higher than for the aPE model, despite having a similar number of parameters (Fig. 3A). Even the base PE(3) model requires more compute than the large aPE model. This enables training aPE models for more epochs under the same compute budget (see Appendix Fig. A4). Scaling the capacity of aPE-L yields substantial gains in precision metrics (Fig. 3B and C). For recall, sample fidelity improves with larger model capacity (as indicated by larger COV-R at small thresholds) but differences start to vanish at $\delta=0.75\mbox{\AA}$ (Fig. 3B). Extrapolation performance also improves for the larger model (see Tab. 4). Similar trends are observed for both base and large variants of PE(3) (see Appendix Fig. A4). This allows our aPE-L model to achieve SOTA results on GEOM-DRUGS for all precision metrics (see Tab. 2). For recall, aPE-L achieves better results than ET-Flow, but worse than DMT-L and MCF-L.
Table 3: Median absolute error of ensemble properties between generated and reference conformers for GEOM-DRUGS. $E$, $\Delta\epsilon$, $E_{\text{min}}$ are in kcal/mol, and $\mu$ is in debye. Best results in bold, second best underlined; our models are marked with an asterisk (*). Results for MCF, ET-Flow, and ours are averaged over three random seeds. See Tab. A11 for results including standard deviations.
Method $E$$\downarrow$ $\mu$$\downarrow$ $\Delta\epsilon$$\downarrow$ $E_{\text{min}}$$\downarrow$
OMEGA 0.68 0.66 0.68 0.69
Torsional Diff. 0.22 0.35 0.54 0.13
MCF-L 0.68 0.28 0.63 0.04
ET-Flow 0.18 0.18 0.35 0.02
*DiTMC+aPE-B 0.17 0.16 0.27 0.01
*DiTMC+aPE-L 0.16 0.14 0.27 0.01
Table 4: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS reporting median RMSD in Å. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Tab. A13 for results including standard deviations.
Method AMR-R $\downarrow$ AMR-P $\downarrow$
102 molecules
MCF - L (242M) 1.60 2.43
*DiTMC+aPE-B 1.60 2.49
*DiTMC+aPE-L 1.51 2.30
77 molecules
MCF - L (242M) 1.51 2.26
*DiTMC+aPE-B 1.47 2.24
*DiTMC+aPE-L 1.28 2.14
@@
Ensemble properties. To complement RMSD-based geometric evaluation with a chemically meaningful assessment, we report the median absolute error (MAE) of ensemble properties between generated and reference conformers. The RMSD metric can penalize conformers that differ due to rotations or atom reordering, but are otherwise chemically equivalent. For our analysis we follow the protocol of MCF [47] and ET-Flow [32] (see Appendix E.3) and compare energy $E$, dipole moment $\mu$, HOMO-LUMO gap $\Delta\epsilon$, and minimum energy $E_{\text{min}}$. Our aPE-L model predicts ensemble properties more accurately than all baselines, highlighting the physical validity of our generated structures. In particular, MCF-L, which shows better performance for recall metrics, is outperformed by a large margin (up to a factor of $4$ for energy).
@@
Generalization performance. We assess the generalization of our model to larger and unseen molecules using the GEOM-XL dataset [46]. It comprises 102 molecules of size $N>100$ atoms, whereas the molecules in the training data contain only $N=44$ atoms on average. Following MCF [47] we report the generalization performance for all 102 molecules and a subset of 77 molecules. Our models perform on par or better to the previously best-performing method MCF-L while using only a fraction of the parameters (Tab. 4). Other baselines (like ET-Flow) are outperformed by a larger margin (see Appendix Tab. A13).

### v2

@@ 4.2 Positional Embeddings
such that $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$. This kind of positional embedding does not preserve rotational nor translational invariance, and serves as a baseline without any symmetry constraints.
Relative Positional Embeddings We use displacements vectors $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ to build pairwise relative Positional Embeddings (rPE) as
$\boldsymbol{p}_{ij}^{\text{rPE}}=\text{MLP}(\vec{r}_{ij})\,$, (9)
such that $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$. This formulation ensures translational invariance but not rotational invariance.
Euclidean Positional Embeddings Adapting ideas from equivariant message passing neural networks like PaiNN [schutt2021equivariant] or NequIP [batzner20223], we construct SO(3)-equivariant pairwise Euclidean Positional Embeddings (PE(3)) as a concatenation of $L+1$ components
@@
where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0,\dots,L$. The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$ and (after concatenation) $\boldsymbol{p}_{ij}^{\text{PE(3)}}\in\mathbb{R}^{(L+1)^{2}\times H}$. Under rotation of the input positions, these positional embeddings transform equivariantly (see Appendix E). Moreover, because displacement vectors are used as inputs, the embeddings are also invariant to translations. As a result, they respect the full set of Euclidean symmetries relevant to molecular geometry.
4.3 DiTMC Block
Based on the positional embedding strategies introduced in Sec. 4.2, we can define different DiTMC blocks that preserve the extent of Euclidean symmetries encoded in the embeddings throughout the model.
Each DiTMC block transforms a set of input tokens $\mathcal{H}$ into a set of output tokens $\mathcal{H}^{\prime}$, which serve as input for the next block. For aPE and rPE, we have $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$, and for PE(3), we have $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}\}$.
In each DiTMC block, we inject time-based, as well as graph-based atom-wise and pair-wise conditioning information via the conditioning tokens introduced in Sec. 4.1. We use the pair-wise graph conditioning tokens $\mathcal{C}_{\text{Pair}}^{\mathcal{G}}$ during the self-attention update,
$\boldsymbol{h}_{i}=\boldsymbol{h}_{i}+\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}_{\text{Pair}}^{\mathcal{G}})_{i}\,$, (11)
where we employ a standard self-attention mechanism for aPE and rPE and an SO(3)-equivariant self-attention mechanism for PE(3). Additionally, we use the time conditioning token $\boldsymbol{c}^{t}$ and the atom-wise graph conditioning tokens $\mathcal{C}_{\text{atom}}^{\mathcal{G}}=\{\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}\,|\,i\in[N]\}$ to obtain per-atom bias and scaling parameters,
$\alpha_{1i},\beta_{1i},\gamma_{1i},\alpha_{2i},\beta_{2i},\gamma_{2i}=\text{MLP}(\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}})\,$, (12)
and apply adaptive layer norm (AdaLN) and adaptive scale (AdaScale) [peebles2023DiT] for conditioning (see Fig. 1A) similar to applications of DiTs in image synthesis.
We provide details on the non-equivariant DiTMC blocks based on aPE and rPE in Appendix D.1, and on the SO(3)-equivariant DiTMC block based on PE(3) in Appendix D.2.
@@ 4.3 DiTMC Block
Table 1: Results on GEOM-QM9 for different generative models (parameter counts in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Appendix Tab. A9 for results including standard deviations.
@@
Figure 2: Analysis of SO(3)-equivariant (PE(3)) and non-equivariant (aPE, rPE) model formulations on GEOM-QM9. (A) Mean Coverage Recall (COV-R) versus root mean square deviation (RMSD) threshold $\delta$ to any reference conformer. (B) Histogram of the minimal RMSD per generated sample. (C) Loss as a function of latent time $t$ relative to PE(3) loss (see Appendix J for details).
5 Experiments
Datasets and Metrics We conduct our experiments on the GEOM dataset [axelrod2022geom], comprising QM9 (133,258 small molecules) and AICures (304,466 drug-like molecules). Reference conformers are generated using CREST [pracht2024crest]. Drug-like molecules exhibit greater structural diversity, including more rotatable bonds and multiple stereocenters. The data splits are taken from Ref. [ganea2021geomol].
We evaluate our models’ ability to generate accurate and diverse conformers using average minimum RMSD (AMR) and coverage (COV), measuring recall (ground-truth coverage) and precision (generation accuracy). A generated conformer is considered valid if it falls within a specified RMSD threshold of any reference conformer ($\delta=0.5\mbox{\AA }$ for GEOM-QM9 and $\delta=0.75\mbox{\AA }$ for GEOM-DRUGS). Following prior work, we generate $2K$ conformers per test molecule with $K$ reference structures. Appendix G.3 provides further details on the calculation of metrics. Following ET-Flow [hassan2024etflow] and GeomMol [ganea2021geomol] we also apply chirality correction (see Appendix G.4).
@@ 5 Experiments
Ablating Self-Attention and Positional Embedding Strategies The modular structure of DiTMC allows efficient exploration of the design space through variations in the positional embeddings and associated attention blocks (see Sec. 4). We define three model variants, DiTMC+aPE, DiTMC+rPE, and DiTMC+PE(3), which differ only in the choice of positional embedding and self-attention formulation. Architectural details can be found in Appendix D. On GEOM-QM9, our models produce diverse, high quality samples, outperforming the current state-of-the-art across all AMR-R, AMR-P, and COV-P metrics, demonstrating the broad applicability of our modular design and conditioning (see Tab. 1). We use the harmonic prior introduced in Ref. [stark2023harmonic] throughout, which yields improved results compared to the Gaussian prior (see Appendix Tab. A13), and therefore adopt it in all subsequent experiments. Notably, our models maintain competitive performance even when using the Gaussian prior, contrary to the findings in Ref. [hassan2024etflow].
@@
Probing the effect of Euclidean symmetries Our PEs form a hierarchy based on the extent of Euclidean symmetry incorporated by construction. This enables a systematic evaluation of how incorporating symmetry affects model behavior. We summarize our findings on GEOM-QM9 below.
@@
Equivariance improves the fidelity of samples. We analyze COV-R Mean as a function of RMSD threshold $\delta$ for all DiTMC variants (see Fig. 2A). The SO(3)-equivariant DiTMC+PE(3) outperforms the non-equivariant DiTMC+aPE and DiTMC+rPE at low $\delta$, indicating that many of the generated conformers closely match the ground-truth structures. This appears as a leftward shift in the distribution of the minimal RMSD found per generated structure (as shown in Fig. 2B) and aligns with the observation that DiTMC+PE(3) achieves better AMR-R and AMR-P values (as reported in Tab. 1). To better understand this behavior, we examine the loss over time $t$ and find that the non-equivariant models exhibit higher error near the data distribution ($t=1$) (see Fig. 2C). The increase in error towards the end of the generation trajectory results in noisier structures and reduced fidelity.
@@
Equivariance increases the computational cost for models of similar size. The benefit of higher fidelity comes at increased computational cost. During training, the equivariant DiTMC+PE(3) is approximately $3.5$ times slower than the non-equivariant DiTMC+aPE and DiTMC+rPE, and about $3$ times slower at inference (see Fig. 3B). All models use the same number of layers and differ only in the number of heads per layer in order to match the total parameter count, as discussed in Appendix D.
Conditioning Strategies To assess the impact of graph conditioning, we compare different conditioning strategies. As further discussed in Appendix I.2, we compare conditioning solely on atom-wise information (Eq. 6) with an extended scheme that also incorporates pair-wise geodesic graph distances (Eq. 7). In Tab. 2, we show ablation results on GEOM-DRUGS. We also ablate conditioning on pairwise information extracted from our conditioning GNN for each edge corresponding to a chemical bond (see Appendix I.2). We find all variants to be effective, but our proposed combination of geodesic distances and atom-wise information to perform best. This experiment underlines the importance of deriving conditioning tokens for all atom pairs, not just those connected by edges in the molecular graph (i.e. by chemical bonds), which lack global information about the graph structure.
@@
Table 2: Ablation of conditioning strategies using DiTMC+aPE-B on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds. See Appendix Tab. A14 for results on GEOM-QM9.
Mean Median Mean Median Mean Median Mean Median
No conditioning 16.6 4.7 1.132 1.110 5.15 1.2 1.849 1.845
Atom-wise 72.8 77.5 0.555 0.565 55.6 53.3 0.762 0.716
Atom-wise & Pair-wise 79.9 85.4 0.434 0.389 76.5 83.6 0.500 0.423
Model Scaling To investigate model scaling, we define a small (“S”), base (“B”), and large (“L”) DiTMC+aPE model variant (see Appendix D). Each model is trained with identical training hyperparameters on the GEOM-DRUGS dataset. We find strong relative improvements for all metrics (up to 54.1% for COV-P Mean) when scaling DiTMC+aPE-S to DiTMC+aPE-B (see Fig. 3A). Scaling DiTMC+aPE-B to DiTMC+aPE-L yields relative improvements, which however might be smaller than expected given the strong increase in performance from the small to the base model. Prior work has shown that to avoid diminishing returns, dataset and model sizes must be scaled together [kaplan2020scaling, hoffmann2022training]. Scaling only the model yields limited benefits as soon as model size saturates given the amount of data. Strong improvements of DiTMC+aPE-L over DiTMC+aPE-B are observed in terms of sampling fidelity (as indicated by higher COV-R and COV-P Mean values at small thresholds $\delta$) but differences start to vanish at $\delta=0.75\mbox{\AA }$ (see Fig. 4A). Moreover, we obtain relative improvement in terms of generalization to larger and previously unseen molecules from the GEOM-XL dataset [jing2022torsional] between 3.1% (for AMR-P Mean) and 7.3% (for AMR-P Median), i.e., DiTMC+aPE-L exhibits significantly stronger out-of-distribution performance compared to the smaller DiTMC+aPE-B (see Tab. 5).
Figure 3: (A) Coverage (COV) mean and Absolute Minimum RMSD (AMR) mean for DiTMC+aPE models of increasing model capacity on GEOM-DRUGS. (B) Training and inference time for different positional embedding (PE) and associated self-attention strategies. (C) Recall and precision AMR mean for the different DiTMC models on GEOM-DRUGS. For panels (B) and (C) we use results from the base (“B”) variant of each model.
Drug-like Molecules For the following experiments, we restrict our analysis to the base (“B”) and large (“L”) DiTMC variants. In contrast to the smaller GEOM-QM9 dataset, the size of GEOM-DRUGS allows us to explore the design space of DiTMC under a more realistic setting, where training is contstrained by a fixed compute budget [hoffmann2022training]. Our budget of 9 GPU days allows training DiTMC+aPE-L and DiTMC+rPE-L for 50 epochs, and DiTMC+PE(3)-L for 10 epochs. For consistency, we train the base models for the same number of epochs as their large counterparts. All DiTMC variants achieve state-of-the-art performance on GEOM-DRUGS for COV-P and AMR-P (see Tab. 3). Importantly, even the smaller DiTMC+aPE-B and DiTMC+PE(3)-B models outperform ET-Flow-SS of similar model size across all metrics, underlining the effectiveness of our approach.
Table 3: Results on GEOM-DRUGS for different generative models (parameter counts in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Appendix Tab. A10 for results including standard deviations.
@@
*DiTMC+aPE-B (9.5M) 79.9 85.4 0.434 0.389 76.5 83.6 0.500 0.423
*DiTMC+rPE-B (9.6M) 79.3 84.6 0.444 0.400 77.2 84.6 0.492 0.414
*DiTMC+PE(3)-B (8.6M) 80.8 85.6 0.427 0.396 75.3 82.0 0.515 0.437
*DiTMC+aPE-L (28.2M) 79.2 84.4 0.432 0.386 77.8 85.7 0.470 0.387
*DiTMC+rPE-L (28.3M) 78.7 84.1 0.438 0.388 78.1 86.4 0.466 0.381
*DiTMC+PE(3)-L (31.1M) 80.8 85.6 0.415 0.376 76.4 82.6 0.491 0.414
Figure 4: (A) Coverage (COV) mean as function of Root Mean Square Deviation (RMSD) threshold $\delta$ and (B) average minimum RMSD (AMR) mean vs. time per conformer for DiTMC+aPE and other state-of-the-art models. Per model markers from left to right correspond to 5, 10, 20, and 50 sampler steps following Refs. [hassan2024etflow, wang2023swallowing]. Note, that the original MCF paper reports results with two different samplers. Benchmark results (Tab. 3) are obtained with DDPM sampler (1000 steps) and AMR vs. time results are reported for DDIM sampler (5–50 steps). (C) Comparison of conformers generated by MCF, ET-Flow, and DiTMC against ground-truth reference conformers from GEOM-XL. Generated conformers are rotationally aligned with their corresponding reference conformer.
Coverage vs. RMSD Threshold. We further analyse the COV-R Mean and COV-P Mean as a function of RMSD threshold $\delta$ for DiTMC+aPE-B and DiTMC+aPE-L (see Fig. 4A). For small thresholds ($\delta<0.4\mbox{\AA }$) both DiTMC+aPE models outperform all other methods for coverage recall and precision. For larger thresholds ($\rho\geq 0.4\mbox{\AA }$) MCF-L starts to outperform aPE models in terms of COV-R Mean. For COV-P Mean, DiTMC+aPE-B and DiTMC+aPE-L perform better than all other methods for all reasonably small thresholds ($\delta<1.2\mbox{\AA }$). In particular for the most relevant regime where thresholds are small, we see a strong benefit due to model scaling. We find similar results for DiTMC+rPE and DiTMC+PE(3) (see Appendix Fig. A9 and Fig. A10).
Pareto Front. Additionally, we investigate the Pareto front of accuracy and computational efficiency, by plotting model accuracy as a function of wall clock time per generated conformer, following Refs. [hassan2024etflow, wang2023swallowing] (see Appendix G.5 for details). As measure for accuracy we consider the average minimum RMSD (AMR), since it is independent of the RMSD threshold. We report results of DiTMC+aPE-B and DiTMC+aPE-L for AMR-R Mean and AMR-P Mean in Fig. 4B. For AMR-P Mean, both DiTMC+aPE models shift the whole Pareto front, yielding higher accuracy at lower computational cost. Even higher accuracies (at the cost of compute time) can be obtained by scaling DiTMC+aPE-B to DiTMC+aPE-L. For recall, both DiTMC+aPE models shift the Pareto front for little compute times, but most accurate results at increased cost are obtained by MCF-L. Similar results are obtained for DiTMC+rPE (see Appendix Fig. A12), but benefits for DiTMC+PE(3) are limited due to high computational cost of equivariant operations (see Appendix Fig. A13).
@@
Ensemble Properties. To complement the RMSD-based geometric evaluation, we report the median absolute error (MAE) of ensemble properties between generated and reference conformers, following the protocol of MCF [wang2023swallowing] and ET-Flow [hassan2024etflow] (see Appendix G.6 for details). Our models predict ensemble properties more accurately than all baselines, highlighting the physical validity of our generated structures (see Tab. 4). In particular, MCF, which shows better performance for recall metrics, is outperformed by a large margin (up to a factor of four for energy). The strong performance of the aPE-B and aPE-L models underlines the potential of achieving high physical validity without any geometric priors, which have been hypothesized to be one of the reasons for ET-Flow outperforming MCF in the ensemble property task [hassan2024etflow].
Table 4: Median absolute error of ensemble properties between generated and reference conformers for GEOM-DRUGS. Best results in bold, second best underlined; our models are marked with an asterisk (*). Results for MCF, ET-Flow, and ours are averaged over three random seeds. See Appendix Tab. A11 for additional results including standard deviations.
MCF-L (242M) 0.68 0.28 0.63 0.04
ET-Flow (8.3M) 0.18 0.18 0.35 0.02
*DiTMC+aPE-B (9.5M) 0.17 0.16 0.27 0.01
*DiTMC+aPE-L (28.2M) 0.16 0.14 0.27 0.01
@@
Generalization Performance Finally, we assess how well our models, trained on GEOM-DRUGS, generalize to larger and previously unseen molecules using the GEOM-XL dataset [jing2022torsional]. It comprises 102 molecules with more than 100 atoms, whereas the molecules in the training set contain only 44 atoms on average. Following MCF [wang2023swallowing] we report results for all 102 molecules and a subset of 77 molecules. ET-Flow uses a slightly different subset of 75 molecules. Our aPE models perform on par with or better than the previously best-performing method MCF-L (see Tab. 5), while using approximately $8$ times and $25$ times fewer parameters for aPE-L and aPE-B, respectively. The other baselines, such as ET-Flow, are outperformed by a larger margin (see Appendix Tab. A12).
Table 5: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Appendix Tab. A12 for additional results including standard deviations.
Method 75 / 77 mols 102 mols
AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
MCF-L (242M) 1.64 1.51 2.57 2.26 1.97 1.60 2.94 2.43
ET-Flow (8.3M) 2.00 1.80 2.96 2.63 2.31 1.93 3.31 2.84
*DiTMC+aPE-B (9.5M) 1.68 1.47 2.59 2.24 1.96 1.60 2.90 2.48
*DiTMC+aPE-L (28.2M) 1.56 1.28 2.47 2.14 1.88 1.51 2.81 2.30